Dividends on Preferred and Common Stock Pecan Theatre Inc. owns and operates movie theaters throughout Florida and Georgia. Pecan Theatre has declared the following annual dividends over a six-year period: Year 1, $32,000; Year 2, $96,000; Year 3, $144,000; Year 4, $176,000; Year 5, $232,000; and Year 6, $288,000. During the entire period ending December 31 of each year, the outstanding stock of the company was composed of 20,000 shares of cumulative, 4% preferred stock, $100 par, and 100,000 shares of common stock, $10 par. Required: 1. Determine the total dividends and the per-share dividends declared on each class of stock for each of the six years. There were no dividends in arrears at the beginning of Year 1. Summarize the data in tabular form. If required, round your answers to two decimal places. If the amount is zero, please enter "0". Preferred Dividends Common Dividends Year Total Dividends Total Per Share Total Per Share Year 1 $ 32,000 $ $ $ $ Year 2 96,000 Year 3 144,000 Year 4 176,000 Year 5 232,000 Year 6 288,000 $ $ 2. Determine the average annual dividend per share for each class of stock for the six-year period. If required, round your answers to two decimal places. Average annual dividend for preferred $ per share Average annual dividend for common $ per share
Answer #1

Solution 1:

Annual Preferred dividend = 20000*$100*4% = $80,000

Preferred Dividends Common Dividends
Year Total Dividends Total Per Share Total Arrears at year end Total Per Share
Year 1 $32,000 $32,000 $1.60 $48,000 $0 $0.00
Year 2 $96,000 $96,000 $4.80 $32,000 $0 $0.00
Year 3 $1,44,000 $1,12,000 $5.60 $0 $32,000 $0.32
Year 4 $1,76,000 $80,000 $4.00 $0 $96,000 $0.96
Year 5 $2,32,000 $80,000 $4.00 $0 $1,52,000 $1.52
Year 6 $2,88,000 $80,000 $4.00 $0 $2,08,000 $2.08
Total $4,80,000 $4,88,000

Solution 2:

Average annual dividend for preferred = ($480000/ 6 years) / 20000 preferred shares = $4 per share

Average annual dividend for Common = ($488000/ 6 years) / 100000 common shares = $0.81 per share